A youth group decide to go bowling. You can rent the number of lanes that you need for $100. The cost per person will drop by $0.50 if they can get ten more people to attend. How many people were in the original group? (Please solve in algebra from?)
Answer by stanbon(75887)

A youth group decide to go bowling. You can rent the number of lanes that you need for $100. The cost per person will drop by $0.50 if they can get ten more people to attend. How many people were in the original group?
------------------
Let number in original group be "x".
Average cost for each was 100/x dollars.
-----------------------
After adding 10 people average cost is 100/(x+10) dollars
------------------------
EQUATION:
old average - new average = (1/2) dollar
100/x - 100/(x+10) = 1/2
Multiply thru by 2x(x+10) to get:
200(x+10) - 200x = x(x+10)
200 = x^2+10x
x^2+10x-200 = 0
(x+20)(x-10) = 0
Positive answer:
x = 10 (# of people in the original group)
=====================
Cheers,
Stan H.